OpenStudy (anonymous):

show that f(x) = r^5 + r^3 has no local extrama

OpenStudy (anonymous):

extrema*

OpenStudy (anonymous):

I think If I am able to show that function is strictly increasing then it should work. A strictly increasing function cuts x axis only one time, I mean only one intercept on x-axis. hmm for x intercept y= 0 or f(x)=0 f(x) = 0 \[x^5 = -x^3 \] Now through inspection we can see that only for x=0 the equation satisfies

OpenStudy (amistre64):

you mean show that it just smooths down

OpenStudy (amistre64):

id say it has no anything since there is no x value to plug in

OpenStudy (amistre64):

f(x) = r^5 + r^3 f'(x) = r' (5r^4 + 3r^2)

OpenStudy (amistre64):

but thats assuming r is a function of x; otherwise f'(x) = 0

OpenStudy (anonymous):

ah okay I see now I took f(x) = x^5 + x^3

OpenStudy (amistre64):

:) yep

OpenStudy (anonymous):

hmm thanks :)

OpenStudy (anonymous):

a local extreme must have second derivative equal to zero and first derivative change signs. here, second derivative is onyl zero for x=0, and there, first derivative is 5x^4+3x^2, which is positive for both x=1 and x=-1 so it is an inflection point and not an extreme.